Find four consective intergers if the sum of the second and fourth is 48.
Answer by mananth(16946)

let the consequtive integers be x,x+1,x+2,x+3
x+1+x+3=48
2x+4 = 48
add -4
2x+4-4=48-4
2x=44
divide by 2
x= 22
..
the numbers are 22,23,24,25